OpenStudy (anonymous):

Medal!!!!!!!!!!!!!! The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

y = 5x − 15 y = 5x + 15 y = 1 over 5x + 21 over 5 y = 1 over 5x − 29 over 5

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@tester97

OpenStudy (anonymous):

hey @ganeshie8

OpenStudy (anonymous):

could you help me?

OpenStudy (anonymous):

I would graph the first line, then graph that same line through (4 ,5) because a parallel line has the same slope. Here's the two lines graphed. See if you can find the slope-intercept equation of a line for the blue line (the one that is parallel and goes through (4,5), this is the one we are trying to figure out)

OpenStudy (anonymous):

yes i see that now what to do?

OpenStudy (anonymous):

dont take my username seriously

OpenStudy (anonymous):

Now you just need to find the equation of a line for the blue line. It's parallel to y = 5x+1, so we know it has the same slope So we know it will be y = 5x - something, and the something will be where it intercepts the y axis, so where does it intercept the y axis?

OpenStudy (anonymous):

so its y=5x-15

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (anonymous):

Yup, you got it! :)

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (anonymous):

can i ask ya another please

OpenStudy (anonymous):

ill make a new question

OpenStudy (anonymous):

Ok. I can try my best to answer it.

OpenStudy (anonymous):

ok thanks